### v1

# JuStRank: Benchmarking LLM Judges for System Ranking

Abstract
Given the rapid progress of generative AI, there is a pressing need to systematically compare and choose between the numerous models and configurations available. The scale and versatility of such evaluations make the use of LLM-based judges a compelling solution for this challenge. Crucially, this approach requires first to validate the quality of the LLM judge itself. Previous work has focused on instance-based assessment of LLM judges, where a judge is evaluated over a set of responses, or response pairs, while being agnostic to their source systems. We argue that this setting overlooks critical factors affecting system-level ranking, such as a judge’s positive or negative bias towards certain systems. To address this gap, we conduct the first large-scale study of LLM judges as system rankers. System scores are generated by aggregating judgment scores over multiple system outputs, and the judge’s quality is assessed by comparing the resulting system ranking to a human-based ranking. Beyond overall judge assessment, our analysis provides a fine-grained characterization of judge behavior, including their decisiveness and bias.
JuStRank: Benchmarking LLM Judges for System Ranking
Ariel Gera, Odellia Boni, Yotam Perlitz, Roy Bar-Haim, Lilach Eden and Asaf Yehudai IBM Research
1 Introduction
The evaluation of Large Language Models (LLMs) is rapidly adopting the LLM-as-a-judge paradigm Zheng et al. (2023), where automatic evaluations with LLMs complement the use of human annotators, or even replace them altogether. LLM-based judges are increasingly relied upon to conclude which models exhibit superior performance, whether novel training and inference approaches are beneficial, and ultimately which LLM configurations offer a better value proposition to users.
Since relying on an inaccurate judge will likely result in sub-optimal decisions, this trend lends an urgency to evaluating the performance of the LLM judges themselves. Indeed, recent works attempt to benchmark judging capabilities, compiling leaderboards of judge performance Lambert et al. (2024); Tan et al. (2024) as well as analyzing their sensitivities and biases Wang et al. (2023); Thakur et al. (2024); Wei et al. (2024); Bavaresco et al. (2024); Feuer et al. (2024); Liu et al. (2024b); Lee et al. (2024a); Xu et al. (2024); Ye et al. (2024).
Figure 1: Instance and system level judges make different calls: An instance-level judge (top) is used to make decisions about the quality of individual responses (which may be produced by different systems). A system-level judge (bottom) is used to make decisions about the overall quality of systems. For clarity, in this illustration, we focus on pairwise decisions.
Figure 2: System-level judge pipeline. Schematic of our data generation pipeline for judge system rankings.
These works all focus on the instance-level performance of judges. A “good” instance-level judge is expected to make a correct judgment about each response, regardless of the system generating it. For example, given a specific pair of responses, the judge may be asked to determine which one is better (Figure 1, top). This approach is very much in line with prevailing paradigms for model alignment (e.g., RLHF, DPO; Lee et al., 2024b) and synthetic data generation Yehudai et al. (2024); these often rely on LLM judges and reward models for making instance-level pairwise decisions on the quality of individual responses.
Although judges are evaluated based on their instance-level performance, very commonly they are actually used for making system-level decisions; namely, to compare and rank different models or different configurations (Figure 1, bottom). Crucially, even very good instance-level capabilities do not guarantee accurate model ranking; and at the same time, mediocre performance on instances could still yield a very accurate overall ranking (Dorner et al., 2024, §2). Thus, the system-level performance of judges – that is, to what degree they can correctly decide between candidate systems, and produce accurate model performance rankings – remains largely an open question. Furthermore, system-level evaluations can unveil an entire range of under-explored judge qualities, such as being biased towards certain models or making un-calibrated model preference judgments.
In this work we aim to address this gap, and characterize the system-level evaluation capabilities and behaviors of LLM-based judges. To this end, we introduce a novel judge benchmark – JuStRank (Judges for System Ranking). JuStRank compares judges by their ability to correctly rank models, based on agreement with a ground-truth model ranking. JuStRank encompasses a collection of $48$ state-of-the-art judges, including both general-purpose LLMs and reward models. Our large-scale benchmark and analysis allow us to investigate the performance and behavior of judges when ranking systems.
Our contributions are as follows:
1. We introduce JuStRank, the first large-scale benchmark of judges for ranking target systems.
2. We quantify the tendency of a judge to exhibit system bias, where some models are judged “unfairly” (§6.2).
3. We reveal an emergent quality of a system-level judge, its decisiveness factor; decisive judges consistently amplify the gap between strong and weak target systems (§6.1).
2 The Gap in Judge Benchmarking
In this section, we outline why existing estimations of judge performance are insufficient to decide which judge is best at choosing between target systems. (Figure 1, bottom).
At present, users looking for a judge for ranking models, will likely choose it according to the available instance-level judge benchmarks. Yet, from a theoretical standpoint instance-level judge performance does not directly correspond to system-level judge performance Dorner et al. (2024).
More specifically, instance-level judge evaluations focus on how many errors the judge makes, and do not address the distribution of these errors across systems.
For system-level judge evaluation, however, the error distribution plays a key role, as judge errors may distribute unevenly across systems, impacting their induced ranking. For example, a judge may exhibit an unjustifiable preference (positive bias) towards responses from a particular system $A$. Thus, this judge will tend to give this system the wrong ranking, even if it makes very few mistakes on responses from other systems (i.e., has an overall high instance-level accuracy). Hence, a more uniform distribution of errors – reflecting less biased judgment – is a desirable quality for system-level judges, and one that may lead to a more accurate ranking.
Drawing on this observation, our goal here is to construct a system-level benchmark for judges. As a benchmark tailored for system-level evaluation, it will enable reliably estimating a judge’s ability to rank systems; moreover, our ranking-oriented analysis can shed light on judge behaviors and biases, as they occur in real-world data.
3 Task Formulation
In this work we study the use of LLM-based judges for determining the relative quality of systems [^1] Henceforth, we will use the term System to refer to a target model or pipeline that performs a task, and Judge for one that is asked to score (or compare) the quality of such systems. Generative LLMs can act as both systems and judges., over a given set of user instructions (prompts).
Formally, we begin with a set of $L$ systems $\mathbf{S}=\{s_{l}\}_{l=1}^{L}$, and $K$ user instructions $\mathbf{I}=\{i_{k}\}_{k=1}^{K}$. Each system produces a response for each such user instruction, denoted as $R=\{r_{k}^{l}\}_{k,l=1,1}^{k,l=K,L}$, such that $s_{l}(i_{k})=r_{k}^{l}$ (see Figure 2).
Judges $\mathbf{J}=\{j_{p}\}_{p=1}^{P}$ map a pair of instruction $i_{k}$, and system response $r_{k}^{l}$ to a scalar score that estimates the quality of the response. Each judge has a specific realization for performing this score mapping [^2] We note that some realizations, such as the comparative realization in §4.2.2, may incorporate a separate set of responses to perform the judgment., of the form: $j_{p}(i_{k},r_{k}^{l})=Score_{k,l}^{p}$. Once a judge $j_{p}$ scores all $K\times L$ responses, we can define a scores matrix $j_{p}(R)\in\mathbb{R}^{K\times L}$ where $j_{p}(R)_{k,l}=Score_{k,l}^{p}$.
In order to quantify system-level quality, we must apply an aggregation method, $a\in A=\{a\colon\mathbb{R}^{K\times L}\longrightarrow\mathbb{R}^{L}\}$. The aggregation method $a$ maps a scores matrix $j_{p}(R)$ to a system-level vector $V^{p,a}\in\mathbb{R}^{L}$ where each entry, $V^{p,a}_{l}$, is a single overall quality score for system $s_{l}$ by judge $j_{p}$. In turn, ordering the systems scores in $V^{p,a}$ induces a ranking over the systems set $\mathbf{S}$.
We test the performance of judge $j_{p}$ as a ranker by checking the correlation between the ranking induced by $V^{p,a}$ and a golden ranking for $\mathbf{S}$.
Judge Model Realization Aggregation Agreement ($\tau$)
with Gold Ranking
Qwen2.5-72B-Instruct Likert Win-Rate .83
URM-LLaMa-3.1-8B Reward Mean .82
GPT-4o-2024-11-20 Anchor Mean .82
Llama-3-1-405b-instruct-fp8 Numeric Mean .81
Mistral-large-instruct-2407 Likert BT .81
GPT-4o-mini-2024-07-18 Numeric Win-Rate .81
ArmoRM-Llama3-8B-v0.1 Reward Mean .80
Llama-3-1-70b-instruct Numeric Win-Rate .80
Skywork-Llama-3.1-8B-v0.2 Reward Mean .79
Llama-3.1-8B-Instruct TokenProbs Mean .78
Table 1: Top 10 judges by ranking performance. Judges are sorted by the Kendall’s Tau correlation between their overall system ranking and the gold ranking from Chatbot Arena (§4.4). For every judge model, only the best-performing realization and aggregation method is shown. For the full results, refer to Appendix Table LABEL:tab:leaderboard_full.
4 Experimental setup
To explore judge performance and behavior, we utilize responses from multiple systems (§4.1) and run reward model judges (§4.2.1) and LLM judges (§4.2.2) over these responses. To obtain system rankings, we experiment with different aggregation methods (§4.3) over the judge scores. Finally, the resulting rankings are compared against a gold system ranking, obtained from a separate dataset (§4.4).
4.1 System Responses Data
We utilize the Arena Hard v0.1 dataset Li et al. (2024) for a diverse set of instructions and system responses. The dataset uses a curated set of $K=500$ challenging instructions, $I$. As of September $2024$, it includes responses from $L=63$ systems, $S$, totaling about $32$K pairs of instructions and their associated system responses, $R$.
4.2 Generating Judgments
For every judge realization, $j_{p}$, we generate a judgment scores matrix, $j_{p}(R)$, over $R$. In total, we examine $48$ judge realizations, yielding a total of $1.5$M individual judge scores ($63$ systems ×$\times~{}500$ instances ×$\times~{}48$ judge realizations).
4.2.1 Reward Models
We run multiple reward models over $R$. While their exact architectures vary, reward models generally produce a scalar quality score for a given pair of an instruction and a system response.
We utilize the following reward models: ArmoRM-Llama3-8B-v0.1 Wang et al. (2024), Eurus-RM-7b Yuan et al. (2024), InternLM2-7b-reward, InternLM2-20b-reward Cai et al. (2024), Skywork-Reward-Llama-3.1-8B-v0.2 Liu et al. (2024a), Llama-3-OffsetBias-RM-8B Park et al. (2024), GRM-Llama3.2-3B-ft Yang et al. (2024), URM-LLaMa-3.1-8B Lou et al. (2024).
4.2.2 LLM Judge Realizations
Unlike dedicated reward models that produce a single score, generative LLMs can be prompted to judge in multiple ways. Thus, for every LLM we examine several judge realizations.
Absolute judgment - Numeric score (Numeric)
The LLM judge is given an instruction and system response, and is asked to provide a quality score for the response between 00 and $100$.
Absolute judgment - Textual score (Likert)
The judge is asked to provide a quality score of the response on a Likert Likert (1932) scale with $5$ labels: [Very Bad, Bad, Mediocre, Good, Very Good]. We then convert the textual judgments to scores in $[1-5]$.
Absolute judgment - Token probablities (TokenProbs)
The task is framed to the judge as a yes/no question: Is this a good response?. We then extract the top log-probabilities for the first generated token, and specifically look at the probabilities for the tokens yes or no. The judgment score $[0.0-1.0]$ is the sum of probabilities for yes divided by the sum of probabilities for yes and no.
Comparative judgment - Anchor model (Anchor)
Here the judgment task is comparative, i.e., the judge is asked to state a preference between two responses rather than an absolute quality judgment of a given response. Conducting paired comparisons between a system and all other systems is unfeasible; thus, we follow Li et al. (2024) and use the responses of GPT-4-0314 as anchors to which the responses of other systems are compared. Given an anchor response and a system response, we ask the judge which one it prefers. The output is then converted to scores in $[-2,+2]$ (where 00 indicates a tie, and $+1$ / $+2$ indicate slight/strong preference for the system response over the anchor response, respectively).
In total, we collect judgments from $10$ LLMs and $4$ realizations [^3] Prompts for all realizations are provided in Appendix G., yielding $40$ LLM judges. We use the following generative LLM judges: Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024), Mixtral-8x22B-Instruct-v0.1, Mistral-Large-Instruct-2407, Llama-3.1-405B-Instruct Dubey et al. (2024), Llama-3.1-70B-Instruct, Llama-3.1-8B-Instruct, Qwen2.5-72B-Instruct, GPT-4o and GPT-4o-mini.
4.3 Aggregations
Given the raw judgment scores of each judge, $j_{p}(R)$, there are multiple ways to construct a ranking of the $63$ target systems. We calculate rankings using Win-rate aggregation, Mean aggregation, Median aggregation, and BT (Bradley-Terry) aggregation. Details are provided in Appendix B.
Figure 3: Comparison to RewardBench. The plot depicts the relative performance of judges present in both JuStRank and RewardBench Lambert et al. (2024). For comparison, we perform Min-Max normalization over the judge performance scores (accuracy for RewardBench, Kendall’s Tau for our results). Results shown are for the BT aggregation method; the LLM judges use the Anchor realization, which is closest to the setting in RewardBench. Plots for the different RewardBench subsets are shown in Appendix Figure 8.
Figure 4: LLM judge realizations. Kendall’s Tau correlations (±$\pm 95\%$ bootstrapping CI) between the system rankings produced by various LLM judge realizations (§4.2.2) and the gold system ranking from Chatbot Arena. The plot depicts results for the BT aggregation method; for the full results, refer to App. Table LABEL:tab:leaderboard_full.
4.4 Gold Ranking - Chatbot Arena Battles
Human preference data from Chatbot Arena Zheng et al. (2023) serve as our ground-truth reference for the relative quality of systems. Chatbot Arena relies on human-annotated “battles” between system responses to produce a system ranking. We use the English Hard Prompts subset [^4] Chatbot Arena Hard Prompts of their data. We chose this subset as its distribution of user instructions has been shown Li et al. (2024) to match that of our system response data (§4.1). We extract the data and ranking following the official code (see Appendix C).
Given a system ranking produced by a judge, we quantify judge performance via the correlation between its ranking and the reference ranking from Chatbot Arena. Simply put, we assume that a ranking given by a good automated judge would have a high agreement with the ranking compiled from human judgments.
5 JuStRank - Judge Performance Results
Table 1 depicts the $10$ top-performing judges on JuStRank, based on their ranking agreement ($\tau$) with the ground-truth human ranking from Chatbot Arena. For each judge model, the best-performing realization and aggregation method is shown.
As seen in the table, there are both LLMs and reward models that reach decent agreement with the gold ranking. Moreover, several $8$B-parameter reward models are on par with much larger LLMs on the task of system ranking. Thus, we see that reward models, which are explicitly trained to make instance-level decisions between pairs of responses, can excel at the system-level ranking task as well.
Note that an identical correlation score with the ground-truth ranking does not indicate that the judges produce the same ranking; rather, each judge has a different pattern of agreement with the ground-truth. Correlations among the judges themselves are shown in App. Fig. 9.
Comparison to Instance-Level Performance
In Figure 3 we compare our system-level judge leaderboard to the instance-level benchmark RewardBench Lambert et al. (2024). The results demonstrate that better instance-level judges are not always better system rankers, highlighting the discrepancy between the two tasks. Thus, JuStRank offers a novel perspective on judge ability. However, there may be additional factors at play as well. For LLM judges, we use a slightly different realization from the comparative prompts used for RewardBench. Moreover, since creators of reward models aim to do well on RewardBench, it is possible that some newer reward models are slightly overfitted to this test distribution.
Figure 5: Predicted pairwise win-rates. Each point represents a win-rate between a pair of systems $WR(s_{a},s_{b})$ (App. E). The x-axis denotes the gold win-rate from Chatbot Arena, and the y-axis denotes the predicted win-rate as derived from the judge scores. The diagonal marks an exact match between the predicted and gold win-rate; the quadrants signify whether the predicted winning system is the same (green) or different (red) from the gold winning system for this pair. Note that every pair is represented twice (e.g., $WR(s_{a},s_{b})=0.2$, $WR(s_{b},s_{a})=0.8$).
5.1 Effects of LLM Realizations
Figure 4 depicts the performance of the LLM judge models by their realization (§4.2.2). The plot demonstrates that the choice of realization has a considerable effect on the system ranking quality; this appears to be nearly as important as the identity of the LLM used. We confirm this finding using statistical variance analysis (Appendix D).
Many works recommend asking LLMs for comparative rather than absolute judgments Zheng et al. (2023). However, in our experiments the comparative realization (Anchor) exhibits lower performance, with the notable exception of GPT-4o. The best realizations overall were Numeric and Likert, where the judge is asked to provide a verbalized quality score. This is in line with findings from Tian et al. (2023), who report better calibration with verbalized LLM confidence scores. The higher performance for both Numeric and Likert realizations – compared to Anchor and TokenProbs – is statistically significant (App. D).
We also note that each realization induces a characteristic distribution of judge scores, $D^{p}$, such that $Score_{k,l}^{p}\sim{D^{p}}$. Notably, the LLM judges tend to produce particular score values more often than others. Refer to Appendix A for more details.
6 Judge Behavior
Next, we explore more fine-grained judge behaviors, beyond the bottom-line system rankings.
To that end, we focus on the judgment task of pairwise system preference, as this is the foundation of system ranking tasks. As in §5, our aim is to gain an understanding of judge performance and characteristics, by comparing judge behavior on pairwise system preference to ground-truth data.
Pairwise Win-Rates
For every judge $j_{p}$, and for every pair of systems ($s_{a}$, $s_{b}$), the win-rate of $s_{a}$, denoted by $WR^{p}(s_{a},s_{b})$, is the number of instances where it received a higher score than $s_{b}$, divided by the number of non-tied instances (cf. Appendix E). Thus, we calculate the pairwise win-rate for each system pair according to each judge. Note that the win-rates are calculated on the scores matrix $j_{p}(R)$, i.e., before applying an aggregation method.
Gold Win-Rates
Similarly, we extract gold pairwise win-rates, $WR^{g}$, from Chatbot Arena (App. C). $59$ systems appear both in our response data (§4.1) and in Chatbot Arena; in total, we have both judge and gold data for $968$ head-to-head comparisons between pairs of systems.
Figure 6: Beta distribution fit of pairwise win-rates. (a): Judge beta fit example. Each point represents the win-rate between a pair of systems, $WR(s_{a},s_{b})$; the curve and $\alpha$ value describe a fit to the beta distribution (App. F). Plots for all judges are in App. Fig. 11. (b): Decisiveness by judge realization. Cell values denote the decisiveness behaviors of different LLM judge realizations, as described by the $\alpha$ value for their win-rate distribution.
6.1 Some Judges Are Particularly Decisive
Figure 5 depicts the relationship between predicted win-rates and gold win-rates for several judges. The quadrants in the figure indicate whether the judge’s pairwise preference decision is aligned with the gold preference. As can be expected, the judge predictions in Figure 5 are often centered around the ground-truth win-rates determined by humans. But strikingly, some judges exhibit unique prediction patterns, yielding win-rates that are consistently closer to the extremes ($0.0$ / $1.0$) compared to the human data. For instance, for pairs with a ground-truth win-rate of $\sim$0.80.80.80.8, we can see that the predicted win-rate in the judgments of Llama-405B (Fig. 5, right) tends to exceed $0.9$. Put simply, when faced with a response from a strong system, the judge is very likely to prefer it over the response of a less capable system, even where human judges are less decisive.
This sigmoidal win-rate prediction pattern resembles behaviors previously described for classifier calibration Silva Filho et al. (2023), where classifiers may exhibit “overconfidence” in their predicted probabilities. [^5] Note, however, that the behavior in our case does not reflect judge probability scores, but rather the empirical ratio of instances where the responses $\{r_{k}^{l}\}_{l=1}^{l=L}$ of a system $k$ are preferred over those of another system. Thus, following Kull et al. (2017), we quantify judges’ decisive (overconfident) behavior by fitting the cumulative beta distribution function to the win-rate prediction plots. This enables describing judge prediction behavior in terms of a single fit value $\alpha=\beta$, where $\alpha\in[0,\infty]$, $\alpha=1$ represents no over- or under-decisiveness, and larger values represent more decisive behavior (refer to Appendix F for details). Figure 6a and App. Fig. 11 depict the beta curve fit for win-rates of various judges.
Figure 6b compares judge realizations in terms of their decisiveness behavior. We see that LLM judges are usually more decisive when directly asked to provide a quality score, and in particular a textual one (Likert); in contrast, the realization that relies on token probabilities (TokenProbs) does not give rise to such a pattern, and can even result in judge “indecision” (i.e., $\alpha<1$).
This pattern can be explained from two directions. First, the human judgments (§4.4) were collected from multiple individuals, who likely have differing preferences; this may introduce some noise that could lead to less extreme win-rates in the gold data. The other factor is the judges, who may rely on certain heuristics to identify responses from strong systems Feuer et al. (2024), leading to more extreme win-rates in the judge data. While the variance between judges (Fig. 6b) supports the latter, we cannot determine this conclusively.
In practical terms, extreme win-rates can be beneficial to users, as they increase the likelihood of a correct system preference decision given a smaller set of responses (see Ashury Tahan et al., 2024).
6.2 Bias Towards Specific Systems
A major concern when using judges for system preference is judge bias – a judge may treat a specific system “unfairly”, by consistently judging its responses too favorably or too harshly.
We define the bias $B_{s_{a}}^{p}$ of judge $j_{p}$ towards system $s_{a}$ by the expectation over the differences between the predicted win-rate and the gold win-rate, over all systems that $s_{a}$ interacts with. Formally, $B_{s_{a}}^{p}=\mathbb{E}_{s_{b}\in S}(WR^{p}(s_{a},s_{b})-WR^{g}(s_{a},s_{b}))$. In other words, if according to $j_{p}$ the win-rates of system $s_{a}$ are (on average) higher than those in the human data, we will say that $j_{p}$ exhibits positive bias towards it; and if they are lower than the ground-truth, $j_{p}$ would be said to exhibit negative bias towards it.
Note that the decisiveness behavior in §6.1 directly entails a general bias pattern in some judges – namely, a positive bias towards strong systems, and a negative bias towards weak ones. Thus, we calculate a decisiveness-corrected bias, ${B^{\prime}_{s_{a}}}^{p}$, where the gold win-rate $WR^{g}$ is replaced by $WR^{g^{\prime}_{p}}$, i.e., the predicted value for the gold win-rate on the beta distribution fit for judge $j_{p}$ (App. F).
Figure 7: System-specific judge biases. The plot depicts win-rate biases of judges towards specific systems, with respect to the ground-truth win-rates from Chatbot Arena (after correction for the beta distribution fit of each judge). This plot portrays select systems with high bias; the full heat map, including all judge realizations and all systems, is shown in App. Fig. 10b.
We observe some consistent trends of system-specific bias that are common across judges. Figure 7 depicts systems for which there is high bias across judges. For instance, most judges exhibit a strong positive bias towards Athene-70B, to the extent that it is often ranked by them as the #1 system. In contrast, GPT-4-0613, which is $27$th in the gold ranking, receives negative bias, resulting in a median rank of $38$ among the judges.
We also ask whether LLM judges exhibit self-bias Xu et al. (2024), i.e., bias towards the system that uses the same underlying LLM. While we find some instances of self-bias, this is not a consistent effect across judge realizations (App. Table 3).
To quantify the overall propensity of a judge for bias, we measure the standard deviation of its bias over all systems, $\delta$ $=\sigma_{s\in{S}}({B^{\prime}}^{p})$. The bias measure for each judge is presented in App. Table 4.
6.3 Characterizing Judge Behaviors
We have shown that beyond their overall ranking capability (§5), judges exhibit distinct traits in their system-level judgments – in particular, they show different levels of decisiveness6.1), and overall propensities for bias6.2). Interestingly, each of these traits (cf. App. Table 4) is correlated to the ranking quality $\tau$, with $r=0.55$ for the $\alpha$ decisiveness measure, and $r=-0.56$ for the bias propensity $\delta$. At the same time, these marked traits are – by design – uncorrelated with each other ($r=-0.07$ between $\alpha$ and $\delta$). Thus, our analyses reveal global system-level judge traits, ones that remain hidden when assessing judges from an instance-level perspective.
7 Related Work
Applying and assessing automatic metrics for system-level evaluation has been studied for decades, in particular for natural language generation tasks Reiter and Belz (2009); Louis and Nenkova (2013); Deutsch et al. (2022). In the context of LLM-based judges, however, system-level evaluation is still under-explored.
For LLM-based judges and reward models, prior works have opted for an instance-level evaluation approach, curating benchmarks of task outputs with ground-truth quality annotations in order to evaluate judge performance. Most prominently, RewardBench Lambert et al. (2024) compares dozens of judges (including reward models, generative LLMs, and classifiers) on the task of correctly deciding between pairs of outputs, labeled as "preferred" or "rejected" by human annotators. RewardBench aims to identify the most suitable judges for model alignment, e.g., for use in RLHF; in contrast, the present work measures judges in terms of their ability to compare the performance of candidate systems. Another recent instance-level benchmark is JudgeBench Tan et al. (2024), which focuses on curating challenging response pairs where the judge must discern subtle errors.
Multiple works are dedicated to analyzing various biases Ye et al. (2024) and undesirable behaviors exhibited by judges. These include positional bias Wang et al. (2023), verbosity bias Saito et al. (2023); Chen et al. (2024) and self-bias Xu et al. (2024), as well as sensitivity to prompts Wei et al. (2024), source datasets Bavaresco et al. (2024), epistemic markers Lee et al. (2024a) and style Feuer et al. (2024); Liu et al. (2024b).
Several popular benchmarks rely on LLM judges to produce leaderboards of state-of-the-art systems. Such benchmarks – e.g., Arena Hard Li et al. (2024) and AlpacaEval Dubois et al. (2024) – do perform a system-level validation of their resulting leaderboards against other benchmark rankings (see Perlitz et al., 2024). However, such efforts are limited to validating the particular dataset and judge setup chosen for the benchmark (usually incorporating GPT-4 as the judge), rather than comparing and analyzing the performance of different judge models and implementations. Thakur et al., 2024 conduct a task-specific system-level evaluation of judges, over the TriviaQA Joshi et al. (2017) dataset. Compared to their work, the present study is on a larger scale and offers novel metrics and analyses on system-level judge behaviors.
8 Discussion
The usage of LLM-based judges is continually expanding. Moreover, many research papers – proposing novel architectures, algorithms and training methods – rely heavily on system-level evaluations using judges as evidence for the utility of their approach. But without evaluating the judges on such system-level tasks, how can one know whether to trust such evaluations, and their conclusions?
We are the first to investigate on a large scale the performance of LLM-based judges on the system ranking task. Our resulting benchmark, JuStRank, will assist users and researchers in choosing the judge best suited for their needs.
Choosing a judge requires many fine-grained decisions. A user can decide which reward model or LLM to use as the judge; opt for relative judgments or absolute scores; try various prompts; apply different aggregations to compile a ranking, etc. Furthermore, these decisions may interact in non-trivial ways (e.g., the distribution of scores a judge tends to output can dictate which aggregations will work well). Indeed, our findings confirm that such decisions substantially affect system-level judgments (§5), and thus are quite likely to change the model selection of an end user, or flip the conclusions of an NLP research paper.
Our system-level approach has multiple additional benefits. First, it forces the evaluation of judges to be representative with respect to the distribution of systems that generate the responses. In existing instance-level benchmarks this factor is not taken into account, and likely results in less accurate judge evaluations.
Second, it affords a new perspective on what it means for a judge to be biased; on the one hand, we discover some decisiveness trends (§6.1) that may actually be useful for making correct preference decisions, and increasing the separability between systems; and on the other, we report some problematic biases that directly distort the judgment of particular systems (§6.2). An important avenue for future work is to connect our findings here to the existing literature on judge biases Ye et al. (2024), and understand to what extent both of these behaviors stem from particular LLM style attributes Feuer et al. (2024).
Given this vast and complex space, our work is admittedly only a first step in understanding the behavior of judges for ranking and selecting LLMs. We encourage the community to explore these issues further: for instance, by training dedicated system-level judges, exploring judge ensembles, or studying other aggregation approaches. We believe that JuStRank can facilitate such research directions, as it can be easily extended to new judges without requiring additional human annotations.
Our hope is that both practitioners and researchers can benefit from JuStRank, by making more informed choices of judges to suit their needs.
9 Conclusion
In this work we conducted the first comprehensive evaluation of system ranking by LLM judges. We tested a wide array of judges, including reward models, as well as different realizations of generative LLMs, over a large collection of systems. We collected system responses over a diverse set of instructions. The judges scored each response, and we compiled a ranking by aggregating the judgments over all the responses. Then, the quality of the judge’s system ranking was compared to a human-based ranking, producing the JuStRank leaderboard.
JuStRank allows users to pick judges that are better aligned with the goal of choosing between different models and configurations. JuStRank demonstrates that judge ranking abilities are not directly tied to LLM size or overall quality, and that some dedicated reward models are on par with leading LLM judges. Moreover, our analysis reveals emergent judge traits – decisiveness and bias – that are strongly correlated with their ranking ability.
Limitations
The gold reference data – the English Hard Prompts subset of Chatbot Arena – does not include user instructions or responses. Hence, we collect judgment data over Arena Hard, which contains a large set of instructions and responses. This raises some questions regarding our ability to directly compare the LLM judges and human judges. However, given that Arena Hard was designed to match the distribution of user instructions in English Hard Prompts (see Li et al., 2024), we assume that these datasets are sufficiently similar.
Our analyses of LLM judge realizations are, by necessity, limited to the specific realization prompts that we used. Several studies show that LLMs Mizrahi et al. (2024) as well as LLM judges Wei et al. (2024) are brittle with respect to prompt phrasing, and hence this may have had an impact on the results.
As in multiple other works, here we treat human preference as a single concept. In practice, however, preference is inherently subjective, and is composed of numerous dimensions (e.g., helpfulness, safety, style, coherence etc.). For instance, one individual may prefer succinct model responses while another would prefer more detailed answers. Thus there is no single “human preference”, but rather a collection of preference decisions that depend on the annotation guidelines, cultural context, and human idiosyncrasies Conitzer et al. (2024); Kirk et al. (2024).
Note that following Peyrard et al. (2021), as well as Chatbot Arena Chiang et al. (2024), we generally regard the ground-truth quality of a system in terms of the Bradley-Terry model; simply put, a better system is a system that “wins” more often. Thus, in this work we do not directly consider the quality difference in system responses per instance, i.e., beyond counting wins/losses. Still, some of the aggregation methods we use (e.g., mean) implicitly reflect other perspectives on system quality.
All of our analyses are performed on heterogeneous datasets of user instructions to LLMs. Thus, while we study judges through the lens of general-purpose LLM usage, we cannot draw conclusions on judge behavior that is task-specific (or in specialized domains), nor on performance in languages other than English Gureja et al. (2024). The issue of task, domain, and language-specific judge behavior is thus an important avenue for future work.
Appendix A Judge Score Distributions
Figure 12 depicts the score distributions, $D^{p}$, of the judges in our data.
Reward model distributions
The reward models exhibit continuous score distributions. As seen in Figure 12, these distributions vary in the range of scores, as well as in the shape of the distribution. Some reward model judges have a narrow range of scores, e.g., $-0.1$ to $0.4$, whereas in others it is much wider, e.g., $-3000$ to $5000$. Similarly, some distributions are more symmetric while others have peaks at more extreme values. However, all distributions are uni-modal, with a single peak. Moreover, we note that the continuous nature of these judgment scores also entails an absence of ties between the judged responses.
LLM Numeric distributions
As shown in Figure 12, even though the LLM judges are given a wide range of possible judgment scores ($[0-100]$), in practice they tend to prefer specific score values. This results in many ties when comparing responses from different systems.
LLM Likert distributions
Similarly to the Numeric distributions, the Likert realizations put most of their probability mass on specific scores, which leads to an even greater inclination towards ties (as here they are limited to a smaller range of scores).
LLM TokenProbs distributions
TokenProbs scores tend to be extreme, namely very close to either $0.0$ or $1.0$. Thus, in many cases the score gap between responses is extremely small. This can result in low judge robustness (see the error bars in Figure 4), as well as a higher sensitivity to the choice of aggregation method.
LLM Anchor distributions
The distribution for Anchor judgments is mainly tied to the quality of the anchor system relative to the other systems. However, we see that it is also affected by the characteristics of the judge. For example, we see in Fig. 12 that Llama-3.1-8B exhibits indecision, rating most responses as comparable to those of the anchor. In addition, for some judges, the proportion of $-1$ scores (i.e., the response is slightly worse than the anchor) or $1$ scores (the response is slightly better than the anchor) is unusually low.
Appendix B Aggregation Methods
Given the raw judgments of each judge, $j_{p}(R)$, there are multiple aggregation methods, $a$, that construct a ranking over all the target systems. Here, we calculate rankings using Win-rate aggregation, BT aggregation, Mean aggregation, and Median aggregation. In the following, we provide further details on each aggregation.
Mean & Median Aggregation
These aggregation methods map a score for each system, $s_{l}$, by relying solely on the scores assigned to its responses by judge $j_{p}$. In other words, the mapping of $V_{l}^{p,a}$ by $a$ depends only on the column corresponding to system $s_{l}$ in $j_{p}(R)$. Accordingly, these aggregations can be viewed as an operation on the columns of the scores matrix $j_{p}(R)$. Specifically, for the Mean aggregation, $V_{l}^{p,a}=\frac{1}{K}\Sigma_{k=1}^{K}Score^{p}_{k,l}$. Similarly, Median aggregation is the median of the vector $j_{p}(R)_{*l}$.
We note that for realizations with discrete score distributions (see §A), many systems will likely share the same median score; in this case, the Median aggregation method fails to separate the systems. Hence, Table LABEL:tab:leaderboard_full contains only a handful of LLM judges with Median aggregation, all using the TokenProbs realization.
Win-rate Aggregation
This aggregation maps each system based on its proportion of wins over other systems, averaged over all instructions $i_{k}\in I$. Formally, $V_{b}^{p,a}=\frac{1}{K}\Sigma_{k=1}^{K}\frac{1}{L-1}\Sigma_{l=1,l\neq b}^{L}\mathbb{I}(Score^{p}_{k,b}>Score^{p}_{k,l})$, where $\mathbb{I}(\cdot)$ denotes the indicator function.
Bradley-Terry Aggregation
Following Chiang et al. (2024), we use the vector of Bradley-Terry (BT) coefficients Bradley and Terry (1952) as system scores.
For calculating the BT scores we use the implementation of the Chatbot Arena official notebook [^6] Arena official notebook. Whereas Chiang et al. (2024) apply this method for battles between responses with a human judge, we apply it over our LLM-based judge data, i.e., each “battle” is a comparison between the judge scores $Score^{p}_{k,a}$, $Score^{p}_{k,b}$ for a response generated by systems $s_{a}$ and $s_{b}$.
When there are no ties, e.g., for the reward model judges, this aggregation produces similar rankings to the win-rate aggregation.
Appendix C Chatbot Arena Data
The data for the Chatbot Arena LLM leaderboard (https://lmarena.ai) consists of "battles" between systems over the same instructions. In these battles, users indicate a preference (or a tie) between a pair of responses generated by different LLMs Zheng et al. (2023); Chiang et al. (2024).
We use their public data file from August 2024777Chatbot Arena data, and follow the official notebook00footnotemark: 0 to extract the raw data, deduplicate it, and calculate the overall system rankings. This dataset includes the human preference judgments and names of the participating systems, but not the instructions or system responses for the battles.
Here we limit the analysis to the English Hard Prompts subset of their data [^8] Chatbot Arena Hard Prompts ($300$K battles). Notably, Arena Hard was specifically designed to match the distribution of user instructions in the English Hard Prompts subset, as described by Li et al. (2024). We follow their code to construct a full system ranking based on these $300$K battles, using Bradley-Terry coefficients. This yields a score for each system in their data, including $59$ systems that are also in our system responses data (§4.1)
Out of this full English Hard data, we also extract a total of $113$K battles that were not judged by humans as ties, and that are between pairs of systems which appear in our responses data. We then use those to calculate win-rates between pairs of systems (§E), yielding a total of $968$ system pairwise win-rates. Note that the Chatbot Arena data does not contain battles between every possible pairing of systems, and thus we do not have win-rates for all combinations of the $59$ systems under consideration. In addition, we limit the analysis to system pairs with at least $10$ non-tied battles.
Appendix D Statistical Analysis of Judge Performance
In §5 and Table LABEL:tab:leaderboard_full we report results of agreement with the gold ranking ($\tau$) for various judge pipelines. Each pipeline consists of a chosen judge model, a realization (§4.2.2) and an aggregation method (§4.3, App. B).
We focus on the LLM judges and perform a three-way ANOVA (analysis of variance), with the ranking correlation $\tau$ as a dependent variable and the model, realization and aggregation as factors. In addition to the variance analysis estimating the effects of these factors, we perform post-hoc pairwise comparisons to ask whether certain configurations (i.e., a specific realization/aggregation) outperform the others. We conduct all analyses using IBM SPSS Statistics v30.0.
The ANOVA shows that both the judge model and the realization have a strong influence on $\tau$, with an effect size (Partial Eta-Squared) of $\eta^{2}=0.81$ for the judge model ($p<0.001$; $F=36.0$), $\eta^{2}=0.51$ for the realization ($p<0.001$; $F=26.6$), and $\eta^{2}=0.78$ for the interaction effect between model and realization ($p<0.001$; $F=10.1$). In contrast, the aggregation methods were not found to have a significant effect on $\tau$ ($\eta^{2}=0.02$; $p>0.5$).
We also perform Tukey’s HSD Tukey (1949) post-hoc tests to compare the means of the variables. The analysis indicates that the both the Numeric (mean $\tau=0.75$; $\sigma_{\tau}=0.06$) and Likert ($\tau=0.74$; $\sigma_{\tau}=0.07$) realizations are significantly better than the Anchor ($\tau=0.71$; $\sigma_{\tau}=0.07$) and TokenProbs ($\tau=0.68$; $\sigma_{\tau}=0.13$) realizations (all $p$ values $<=0.002$). The differences between aggregation methods are not statistically significant.
Appendix E Pairwise Win-Rates
We denote the win-rate of system $s_{a}$ over system $s_{b}$ as $WR(s_{a},s_{b})^{p}$ where $p$ denotes the judge upon which the win-rate was calculated, and $p\in J\cup\{g\}$, where $g$ stands for human gold data.
The win-rate of system $s_{a}$ over system $s_{b}$ according to judge $j_{p}$ over the set of instances $I$ is calculated as the proportion of instances where the score given by $j_{p}$ to the response generated by $s_{a}$ surpasses that of system $s_{b}$, where ties are excluded. Namely $WR^{p}(s_{a},s_{b})=\frac{1}{K-|T^{p}_{a,b}|}\Sigma_{k=1}^{K}\mathbb{I}(Score^{p}_{k,a}>Score^{p}_{k,b})$ Where $T^{p}_{a,b}=\{i_{k}|Score^{p}_{k,a}=Score^{p}_{k,b}\}$, and $\mathbb{I}(\cdot)$ denotes the indicator function. Notice that $WR^{p}(s_{a},s_{b})=1-WR^{p}(s_{b},s_{a})$.
To quantify the agreement between the judge and gold win-rates we also define an Accuracy metric. This measures the proportion of pairs where the judge pairwise system preference decisions are in agreement with those of the human gold-data. In other words, we want to count the pairs that appear in the first and third quadrants in Figure 5; namely, the pairs where the judge and gold win-rate are both bigger than $0.5$, or the pairs where both are lower than $0.5$, representing agreement on the winning system. For that, we denote all the pairs of systems we have in the gold data as $\{s_{a^{m}},s_{b^{m}}\}_{m=1}^{M}$. Now the Accuracy is defined as follows:
$Acc_{WR}^{p}=\frac{1}{M}\Sigma_{m=1}^{M}\mathbb{I}(\mathbb{I}(WR^{p}(s_{a^{m}},s_{b^{m}})>0.5)$
$=\mathbb{I}(WR^{g}(s_{a^{m}},s_{b^{m}})>0.5))$
Additionally, we define a second metric, the Mean Squared Error over all win-rate pairs.
$MSE_{WR}^{m}=\frac{1}{M}\Sigma_{m=1}^{M}(WR^{g}(s_{a^{m}},s_{b^{m}})$
$-WR^{p}(s_{a^{m}},s_{b^{m}}))^{2}.$
The $Acc_{WR}^{p}$ scores are in high agreement with the JuStRank judge ranking quality scores $\tau$ (Pearson correlation of $r=0.96$ for the BT aggregation, $r=0.79$ for the Mean aggregation). This highlights the direct link between judges’ ability to rank systems and their performance on pairwise system preference.
The $MSE_{WR}^{p}$ scores have a low correlation with the JuStRank judge $\tau$ scores ($r=-0.19$ for the BT aggregation, $r=-0.07$ for the Mean aggregation). This can be explained by the decisiveness effect (§6.1), where judges deviate substantially from the gold win-rate, but mostly toward the stronger system in the pair.
Appendix F Beta Distribution Fit
Following Kull et al. (2017), we model the relation between judge and gold win-rates using the cumulative distribution function (CDF) of the Beta distribution. We parameterize the distribution such that both shape parameters $\alpha$ and $\beta$ are equal ($\alpha=\beta$).
The CDF of the Beta distribution, defined over the interval $[0,1]$, for $\alpha=\beta\in[0,\infty]$ provides a wide range of function fits: a linear $y=x$ fit for $\alpha=1$, a sigmoidal fit for larger $\alpha$ values, and approaching a step function as $\alpha\to\infty$. These attributes make it particularly suited for our data characteristics.
Given a set of data points $\{(WR^{p}(s_{a^{m}},s_{b^{m}}),WR^{g}(s_{a^{m}},s_{b^{m}})\}_{m=1}^{M}$, where $WR^{p}(s_{a^{m}},s_{b^{m}})\in[0,1]$ represents the judge win-rate and $WR^{g}(s_{a^{m}},s_{b^{m}})\in[0,1]$ denotes the gold win-rate between system, $s_{a^{m}}$ and $s_{b^{m}}$. We fit the Beta CDF by optimizing the shape parameter $\alpha$. The optimization objective is minimizing the sum of absolute errors (SAE) between the judge win-rate, $WR^{p}(s_{a^{m}},s_{b^{m}})$, and the predicted values from the Beta CDF. In order to capture the behavior across the entire range of win-rates, we weight the errors by the distance of $WR^{p}$ from $0.5$:
$\text{SAE}=\sum_{m=1}^{M}\gamma(WR^{p}(s_{a^{m}},s_{b^{m}}))\cdot\bigg{|}WR^{p}(s_{a^{m}},s_{b^{m}})$
$-F_{\text{Beta}}(WR^{g}(s_{a^{m}},s_{b^{m}});\alpha)\bigg{|}$
where $F_{\text{Beta}}(x;\alpha)$ denotes the Beta CDF with shape parameters $\alpha=\beta$, and $\gamma$ is the distance of $WR^{p}$ from $0.5$.
The optimization was performed using the scipy.optimize.minimize [^9] SciPy Documentation for scipy.optimize.minimize function, with the parameter ($\alpha$) constrained to a reasonable range $[0.1,10000]$. This approach efficiently identified the best-fit parameter ($\alpha$).
The resulting Beta CDF closely captures the empirical data distribution, as validated both quantitatively, through low SAE, and qualitatively via visual inspection. Figure 11 depicts the fitted Beta CDF curve and the observed data points, demonstrating the effectiveness of this approach for modeling the judges’ predicted win-rate distribution.
Appendix G LLM Judge Prompts
Below we list the prompts we use for each LLM judge realization (§4.2.2).
[h!] Numeric Here is a user input and a model response. On a scale of 00 to $100$, to what extent is this a good response for the given input? Reply with your rating score without any preceding explanation. Input: [user instruction] Response: [system response] Rating (00-$100$):
[h!] Likert Here is a user input and a model response. To what extent is this a good response for the given input? Provide a rating from one of the following choices: ’Very Bad’, ’Bad’, ’Mediocre’, ’Good’, ’Very Good’. Reply using the format of [[rating]], for example: ’[[Mediocre]]’ Input: [user instruction] Response: [system response] Rating:
[ht] TokenProbs Here is a user input and a model response. Is this a good response for the given input? Answer with only yes/no. Input: [user instruction] Response: [system response] Good response? (Yes/No):
[h] Anchor Here is a user input and responses from two assistants, A and B. Which response is better? You must output only one of the following choices as your final verdict with a label: 1. Assistant A is significantly better: [[A>>B]] 2. Assistant A is slightly better: [[A>B]] 3. Tie, relatively the same: [[A=B]] 4. Assistant B is slightly better: [[B>A]] 5. Assistant B is significantly better: [[B>>A] Example output: "My final verdict is tie: [[A=B]]". <|User Prompt|> [user instruction] <|The Start of Assistant A’s Answer|> [system response] <|The End of Assistant A’s Answer|> <|The Start of Assistant B’s Answer|> [anchor system response] <|The End of Assistant B’s Answer|> Final Verdict:
Table 2: Judges by ranking performance. The judges are sorted by the Kendall’s Tau correlation between their overall system ranking and the gold ranking from Chatbot Arena (§4.4).
Judge Model Realization Aggregation Agreement ($\tau$)
w/ Gold Ranking
Qwen2.5-72B-Instruct Likert Win-Rate .827
URM-LLaMa-3.1-8B Reward Mean .823
GPT-4o-2024-11-20 Anchor Mean .822
URM-LLaMa-3.1-8B Reward BT .819
Qwen2.5-72B-Instruct Likert BT .817
URM-LLaMa-3.1-8B Reward Win-Rate .816
Qwen2.5-72B-Instruct Numeric BT .814
GPT-4o-2024-11-20 Anchor Win-Rate .814
Qwen2.5-72B-Instruct Numeric Win-Rate .813
Llama-3-1-405b-instruct-fp8 Numeric Mean .812
Llama-3-1-405b-instruct-fp8 Numeric Win-Rate .812
Mistral-large-instruct-2407 Likert BT .811
GPT-4o-2024-11-20 Anchor BT .809
Mistral-large-instruct-2407 Numeric BT .809
URM-LLaMa-3.1-8B Reward Median .809
GPT-4o-mini-2024-07-18 Numeric Win-Rate .807
Llama-3-1-405b-instruct-fp8 Numeric BT .805
GPT-4o-mini-2024-07-18 Numeric BT .804
Mistral-large-instruct-2407 Numeric Win-Rate .802
Qwen2.5-72B-Instruct Likert Mean .801
ArmoRM-Llama3-8B-v0.1 Reward Mean .800
Qwen2.5-72B-Instruct Anchor Mean .799
GPT-4o-mini-2024-07-18 Likert BT .798
Llama-3-1-70b-instruct Numeric Win-Rate .798
Llama-3-1-70b-instruct Numeric BT .798
Mistral-large-instruct-2407 Likert Win-Rate .798
Qwen2.5-72B-Instruct Anchor BT .794
Llama-3-1-405b-instruct-fp8 Likert Win-Rate .793
Llama-3-1-70b-instruct TokenProbs Win-Rate .793
GPT-4o-mini-2024-07-18 Likert Win-Rate .793
ArmoRM-Llama3-8B-v0.1 Reward Median .793
Llama-3-1-405b-instruct-fp8 Likert BT .787
Mistral-large-instruct-2407 Anchor Win-Rate .786
Skywork-Llama-3.1-8B-v0.2 Reward Mean .786
Qwen2.5-72B-Instruct Anchor Win-Rate .786
Mistral-large-instruct-2407 Likert Mean .782
GPT-4o-mini-2024-07-18 Numeric Mean .781
Skywork-Llama-3.1-8B-v0.2 Reward Win-Rate .780
Llama-3-1-405b-instruct-fp8 Likert Mean .780
Skywork-Llama-3.1-8B-v0.2 Reward BT .778
Llama-3.1-8B-Instruct TokenProbs Mean .778
Qwen2.5-72B-Instruct TokenProbs BT .777
Llama-3.1-8B-Instruct TokenProbs Median .776
Mixtral-8x22B-instruct-v0.1 Numeric BT .776
Llama-3-1-70b-instruct TokenProbs Median .776
GPT-4o-2024-11-20 Numeric BT .774
GPT-4o-mini-2024-07-18 Likert Mean .773
Qwen2.5-72B-Instruct Numeric Mean .773
GPT-4o-2024-11-20 Likert BT .773
GPT-4o-2024-11-20 Numeric Win-Rate .771
Llama-3-OffsetBias-RM-8B Reward Win-Rate .765
Llama-3-1-70b-instruct TokenProbs BT .765
Llama-3-OffsetBias-RM-8B Reward BT .765
Skywork-Llama-3.1-8B-v0.2 Reward Median .764
Llama-3-1-70b-instruct TokenProbs Mean .764
Mistral-large-instruct-2407 Anchor Mean .764
Llama-3-1-70b-instruct Numeric Mean .764
ArmoRM-Llama3-8B-v0.1 Reward BT .763
ArmoRM-Llama3-8B-v0.1 Reward Win-Rate .762
Llama-3-OffsetBias-RM-8B Reward Median .759
GPT-4o-mini-2024-07-18 TokenProbs Win-Rate .759
GPT-4o-2024-11-20 Likert Win-Rate .758
Llama-3-OffsetBias-RM-8B Reward Mean .757
Mixtral-8x22B-instruct-v0.1 Numeric Win-Rate .756
GPT-4o-mini-2024-07-18 TokenProbs BT .752
Qwen2.5-72B-Instruct TokenProbs Median .752
Mistral-large-instruct-2407 Numeric Mean .750
Llama-3-70b-instruct Numeric BT .749
Qwen2.5-72B-Instruct TokenProbs Win-Rate .748
Llama-3-1-405b-instruct-fp8 Anchor Win-Rate .748
Llama-3-1-70b-instruct Likert Mean .746
GPT-4o-2024-11-20 Likert Mean .744
Llama-3.1-8B-Instruct TokenProbs Win-Rate .744
Llama-3-1-405b-instruct-fp8 Anchor Mean .744
Llama-3.1-8B-Instruct TokenProbs BT .741
Llama-3-1-405b-instruct-fp8 TokenProbs Win-Rate .741
GPT-4o-mini-2024-07-18 TokenProbs Mean .741
Mixtral-8x22B-instruct-v0.1 Likert BT .738
GPT-4o-2024-11-20 Numeric Mean .738
Llama-3-1-405b-instruct-fp8 TokenProbs Median .737
Llama-3.1-8B-Instruct Likert Mean .736
Llama-3-70b-instruct Numeric Win-Rate .733
Llama-3-1-405b-instruct-fp8 TokenProbs Mean .733
Llama-3-1-70b-instruct Likert Win-Rate .732
Mixtral-8x22B-instruct-v0.1 Likert Win-Rate .732
Qwen2.5-72B-Instruct TokenProbs Mean .732
Internlm2-7b-reward Reward Mean .731
Llama-3-1-405b-instruct-fp8 Anchor BT .730
Mistral-large-instruct-2407 TokenProbs Mean .730
Internlm2-20b-reward Reward Mean .728
Mistral-large-instruct-2407 Anchor BT .725
Internlm2-20b-reward Reward Median .724
GPT-4o-mini-2024-07-18 TokenProbs Median .723
Llama-3.1-8B-Instruct Likert BT .723
Llama-3-1-70b-instruct Likert BT .722
Internlm2-7b-reward Reward Median .721
Mixtral-8x22B-instruct-v0.1 Likert Mean .719
Internlm2-7b-reward Reward Win-Rate .717
Internlm2-20b-reward Reward BT .717
Mixtral-8x22B-instruct-v0.1 TokenProbs Win-Rate .717
Llama-3-1-70b-instruct Anchor Win-Rate .716
GRM-Llama3.2-3B Reward Mean .716
Internlm2-20b-reward Reward Win-Rate .716
Mixtral-8x22B-instruct-v0.1 Numeric Mean .715
Llama-3-1-70b-instruct Anchor Mean .714
GRM-Llama3.2-3B Reward Win-Rate .712
Internlm2-7b-reward Reward BT .712
GRM-Llama3.2-3B Reward BT .711
GRM-Llama3.2-3B Reward Median .706
GPT-4o-2024-11-20 TokenProbs Median .704
Llama-3-70b-instruct Numeric Mean .704
Mixtral-8x22B-instruct-v0.1 TokenProbs BT .702
GPT-4o-2024-11-20 TokenProbs Mean .701
GPT-4o-2024-11-20 TokenProbs BT .700
Llama-3-70b-instruct Likert BT .698
Llama-3-70b-instruct TokenProbs Win-Rate .696
GPT-4o-2024-11-20 TokenProbs Win-Rate .696
Llama-3.1-8B-Instruct Anchor Mean .695
Llama-3.1-8B-Instruct Likert Win-Rate .694
Llama-3-1-70b-instruct Anchor BT .688
Llama-3-70b-instruct Likert Win-Rate .681
Llama-3.1-8B-Instruct Numeric Mean .680
Llama-3-70b-instruct Likert Mean .678
Llama-3.1-8B-Instruct Anchor BT .677
GPT-4o-mini-2024-07-18 Anchor Mean .675
Llama-3-1-405b-instruct-fp8 TokenProbs BT .672
Llama-3.1-8B-Instruct Numeric BT .668
GPT-4o-mini-2024-07-18 Anchor Win-Rate .668
Llama-3-70b-instruct Anchor Mean .667
Llama-3-70b-instruct TokenProbs Mean .666
Mixtral-8x22B-instruct-v0.1 Anchor Mean .665
Llama-3-70b-instruct TokenProbs BT .663
GPT-4o-mini-2024-07-18 Anchor BT .659
Mixtral-8x7B-instruct-v0.1 Numeric BT .656
Mixtral-8x7B-instruct-v0.1 Anchor BT .655
Mixtral-8x22B-instruct-v0.1 TokenProbs Mean .650
Eurus-RM-7b Reward Median .643
Eurus-RM-7b Reward Mean .641
Mixtral-8x22B-instruct-v0.1 Anchor BT .641
Llama-3.1-8B-Instruct Anchor Win-Rate .639
Llama-3-70b-instruct Anchor Win-Rate .638
Llama-3-70b-instruct Anchor BT .633
Llama-3.1-8B-Instruct Numeric Win-Rate .632
Eurus-RM-7b Reward Win-Rate .629
Eurus-RM-7b Reward BT .628
Mixtral-8x7B-instruct-v0.1 Numeric Win-Rate .626
Mixtral-8x7B-instruct-v0.1 Numeric Mean .626
Mixtral-8x7B-instruct-v0.1 Anchor Win-Rate .622
Mixtral-8x22B-instruct-v0.1 Anchor Win-Rate .612
Mixtral-8x7B-instruct-v0.1 Anchor Mean .610
Mixtral-8x7B-instruct-v0.1 Likert BT .590
Mixtral-8x7B-instruct-v0.1 Likert Mean .585
Mixtral-8x7B-instruct-v0.1 Likert Win-Rate .543
Mixtral-8x7B-instruct-v0.1 TokenProbs BT .427
Mistral-large-instruct-2407 TokenProbs Win-Rate .417
Mixtral-8x7B-instruct-v0.1 TokenProbs Mean .411
Mixtral-8x7B-instruct-v0.1 TokenProbs Win-Rate .371
Mistral-large-instruct-2407 TokenProbs BT .369
Mistral-large-instruct-2407 TokenProbs Median .363
Figure 8: Comparison to RewardBench. The plot depicts the relative performance of judges present in both JuStRank and RewardBench Lambert et al. (2024). For comparison, we perform Min-Max normalization over the judge performance scores (accuracy for RewardBench, Kendall’s Tau for our results). The results shown are for the BT aggregation method; the LLM judges use the Anchor realization, which is closest to the setting in RewardBench. Each panel portrays a different subset of RewardBench.
Figure 9: Judge Correlations. Kendall’s Tau correlations between the system rankings produced by the different judge realizations, using the BT aggregation method. The first row/column denotes correlations with the reference ranking from Chatbot Arena.
Figure 10: System-specific judge biases. The heat maps depict the win-rate biases of various judges towards specific systems (§6.2), with respect to the ground-truth win-rates from Chatbot Arena. (a): Bias w.r.t. the raw ground-truth win-rates $WR^{g}$; (b): Bias w.r.t. the fit value for the gold win-rate $WR^{g^{\prime}}$ on the beta distribution fit (App. F) for each judge.
Figure 11: Beta distribution fit of pairwise win-rates (Part 1/4)
Figure 11: Beta distribution fit of pairwise win-rates (Part 2/4)
Figure 11: Beta distribution fit of pairwise win-rates (Part 3/4)
Figure 11: Beta distribution fit of pairwise win-rates (Part 4/4). Each point represents the win-rate between a pair of systems, $WR(s_{a},s_{b})$; the curve and $\alpha$ value describe a fit to the beta probability distribution. Refer to Appendix F for details.
Judge Self-bias Significance $p$-value
GPT-4o-mini-2024-07-18 (Anchor) $-0.05$
GPT-4o-mini-2024-07-18 (Likert) $-0.04$
GPT-4o-mini-2024-07-18 (Numeric) $+0.03$ $>0.5$ (N.S.)
GPT-4o-mini-2024-07-18 (TokenProbs) $+0.06$ $0.13$ (N.S.)
Llama-3-1-70b-instruct (Anchor) $-0.05$
Llama-3-1-70b-instruct (Likert) $+0.16$ $037.1e-$ 03
Llama-3-1-70b-instruct (Numeric) $-0.00$
Llama-3-1-70b-instruct (TokenProbs) $-0.03$
Llama-3-70b-instruct (Anchor) $+0.09$ $044.7e-$ 04
Llama-3-70b-instruct (Likert) $+0.15$ $088.4e-$ 08
Llama-3-70b-instruct (Numeric) $+0.14$ $131.8e-$ 13
Llama-3-70b-instruct (TokenProbs) $-0.01$
Llama-3.1-8B-Instruct (Anchor) $-0.07$
Llama-3.1-8B-Instruct (Likert) $-0.04$
Llama-3.1-8B-Instruct (Numeric) $+0.02$ $>0.5$ (N.S.)
Llama-3.1-8B-Instruct (TokenProbs) $-0.04$
Mistral-large-instruct-2407 (Anchor) $-0.07$
Mistral-large-instruct-2407 (Likert) $+0.02$ $>0.5$ (N.S.)
Mistral-large-instruct-2407 (Numeric) $+0.06$ $0.33$ (N.S.)
Mistral-large-instruct-2407 (TokenProbs) $+0.01$ $>0.5$ (N.S.)
Table 3: Judge self-bias. The table shows the self-bias values for LLM judge realizations, i.e., the value of the corrected bias ${B^{\prime}_{s_{a}}}^{p}$6.2) where the LLM judge $p$ and system $s_{a}$ correspond to the same underlying LLM. For positive self-bias values we test the statistical significance using paired t-tests (one-sided, with Bonferroni correction). N.S.: non-significant ($p>0.05$).
Figure 12: Judge score distributions (Part 1/3)
Figure 12: Judge score distributions (Part 2/3)
Figure 12: Judge score distributions (Part 3/3).
Judge Model Realization Agreement Decisiveness Bias
with Gold $\tau\uparrow$ $\alpha\uparrow$ $\delta$$\downarrow$
URM-LLaMa-3.1-8B Reward .819 1.84 .085
Qwen2.5-72B-Instruct Likert .817 4.76 .079
Qwen2.5-72B-Instruct Numeric .814 4.09 .079
Mistral-large-instruct-2407 Likert .811 5.47 .086
GPT-4o-2024-11-20 Anchor .809 3.07 .085
Mistral-large-instruct-2407 Numeric .809 3.01 .082
Llama-3-1-405b-instruct-fp8 Numeric .805 4.33 .087
GPT-4o-mini-2024-07-18 Numeric .804 2.91 .077
GPT-4o-mini-2024-07-18 Likert .798 4.61 .087
Llama-3-1-70b-instruct Numeric .798 2.69 .087
Qwen2.5-72B-Instruct Anchor .794 2.93 .090
Llama-3-1-405b-instruct-fp8 Likert .787 5.22 .097
Skywork-Llama-3.1-8B-v0.2 Reward .778 2.46 .100
Qwen2.5-72B-Instruct TokenProbs .777 2.69 .082
Mixtral-8x22B-instruct-v0.1 Numeric .776 2.12 .089
GPT-4o-2024-11-20 Numeric .774 2.15 .077
GPT-4o-2024-11-20 Likert .773 5.49 .089
Llama-3-1-70b-instruct TokenProbs .765 1.26 .070
Llama-3-OffsetBias-RM-8B Reward .765 1.39 .076
ArmoRM-Llama3-8B-v0.1 Reward .763 1.84 .092
GPT-4o-mini-2024-07-18 TokenProbs .752 2.10 .084
Llama-3-70b-instruct Numeric .749 1.27 .084
Llama-3.1-8B-Instruct TokenProbs .741 .598 .061
Mixtral-8x22B-instruct-v0.1 Likert .738 2.53 .108
Llama-3-1-405b-instruct-fp8 Anchor .730 3.58 .112
Mistral-large-instruct-2407 Anchor .725 2.13 .111
Llama-3.1-8B-Instruct Likert .723 .935 .090
Llama-3-1-70b-instruct Likert .722 3.90 .120
Internlm2-20b-reward Reward .717 1.90 .098
Internlm2-7b-reward Reward .712 2.35 .113
GRM-Llama3.2-3B Reward .711 2.30 .114
Mixtral-8x22B-instruct-v0.1 TokenProbs .702 1.85 .088
GPT-4o-2024-11-20 TokenProbs .700 2.22 .093
Llama-3-70b-instruct Likert .698 2.40 .122
Llama-3-1-70b-instruct Anchor .688 2.71 .126
Llama-3.1-8B-Instruct Anchor .677 .868 .085
Llama-3-1-405b-instruct-fp8 TokenProbs .672 1.55 .092
Llama-3.1-8B-Instruct Numeric .668 1.20 .104
Llama-3-70b-instruct TokenProbs .663 .775 .071
GPT-4o-mini-2024-07-18 Anchor .659 1.41 .111
Mixtral-8x7B-instruct-v0.1 Numeric .656 1.27 .102
Mixtral-8x7B-instruct-v0.1 Anchor .655 1.17 .102
Mixtral-8x22B-instruct-v0.1 Anchor .641 1.50 .140
Llama-3-70b-instruct Anchor .633 1.82 .132
Eurus-RM-7b Reward .628 2.49 .138
Mixtral-8x7B-instruct-v0.1 Likert .590 .838 .110
Mixtral-8x7B-instruct-v0.1 TokenProbs .427 .739 .107
Mistral-large-instruct-2407 TokenProbs .369 1.17 .123
Table 4: Judge characteristics. The table presents three measures for each judge realization: an overall ranking quality $\tau$ 5, Kendall’s Tau correlation with the Chatbot Arena gold ranking), a decisiveness score $\alpha$ 6.1, App. F), and its propensity for system-specific biases $\delta$ 6.2). Correlations $\tau$ shown are for the BT aggregation method; $\alpha$ and $\delta$ are calculated on the judge scores before aggregation. $\downarrow$: Lower is better.